Sam Long anticipates he will need approximately $227,400 in 15 years to cover his 3-year-old daughter’s college bills for a 4-year degree.

How much would he have to invest today at an interest rate of 6 percent compounded semiannually? (Do not round intermediate calculations. Round your answer to the nearest cent.)
Answer #1

We use the formula:
A=P(1+r/2)^2n
where
A=future value
P=present value
r=rate of interest
n=time period.

227,400=P*(1+0.06/2)^(2*15)

P=227,400/(1+0.06/2)^(2*15)

=227,400*0.41198676

=$93685.79(Approx).